What is KG equation
1
Expert's answer
2020-12-10T11:10:51-0500

Answer

The Klein–Gordon equation is a relativistic wave equation, related to the Schrödinger equation. It is second-order in space and time and manifestly Lorentz-covariant.

As below

"\nabla^2\psi-\frac{1}{c^2}\frac{ d^2\psi}{dt^2}=\frac{ m^2c^2\psi}{\hbar^2}"